# A forest of gravitational waves in our Galactic Centre

### v1

@@ III.1 XMRI forest
[…]
@@
The results of our XMRI forest analysis are presented in Figure 4, along with several representative gravitational wave sources expected in the LISA band. Representative signals from massive black hole binaries , Galactic verification binaries, and characteristic EMRI harmonics are overlaid for comparison, together with the LISA instrumental sensitivity curve, consistent with those shown in Figure (1) of [1]. The yellow dots represent the superposition of individual XMRI strains binned in frequency, while the blue curve traces the non-zero strain values, illustrating the overall spectral shape of the forest. Notably, the XMRI background partially overlaps with the characteristic harmonics of EMRI signals around frequencies of $\sim 1\,\mathrm{mHz}$, potentially introducing additional confusion noise in this region. Nevertheless, due to the low companion masses in XMRI systems, the background level remains relatively low and is not expected to substantially impede the detection of primary sources such as MBHBs or Galactic binaries.
@@ III.2 E-EMRI forest
Figure 7 shows the forest contribution from E-EMRI sources in our mixed catalog. These sources exhibit high strain amplitudes, producing a pronounced forest in the LISA band. The visible harmonic structure arises from their eccentric orbits, which generate gravitational radiation bursts across multiple harmonics. For the 10-year observation (lower panel), the improved frequency resolution leads to a finer and smoother spectral structure, particularly at higher frequencies ($f\gtrsim 1\,\mathrm{mHz}$). Compared to the 1-year case, the extended observation reduces confusion at low frequencies and sharpens the background contribution at high frequencies.
@@
The distribution of individual EMRI source strain amplitudes (grey dots) exhibits a distinct bifurcation, forming two visibly separated branches. This separation arises from the variation in source SNR. The lower branch is predominantly composed of sources with intermediate SNRs, whereas the upper branch originates from a small subset of loud sources. Owing to our background construction strategy—which maximizes spectral coverage by selecting realizations with the densest frequency occupation—the upper envelope is shaped by cases with high SNRs. These high SNR sources contribute to the strain spectrum at high frequencies, leading to a visibly elevated branch relative to the bulk of the population.
@@ III.3 Extragalactic forests: The Andromeda galaxy
Figure 8 shows characteristic strain spectra for both E-EMRI and XMRI populations from M32 over a 1-year observation period with a handful of sources, as an illustration. The upper panel demonstrates that despite the substantial distance, many E-EMRI signals remain above LISA’s sensitivity curve in the $10^{-4}$ to $10^{-3}$ Hz band, particularly those involving higher-mass compact objects on highly eccentric orbits. In contrast, the lower panel shows that standard XMRI signals experience more severe attenuation, with most falling below the detection threshold across the frequency spectrum. Obviously, understanding the shape of the forest when taking into account all close-by galaxies is important, and this will be addressed in a separate work.
@@ IV Conclusions
Acknowledgments
We thank Leor Barack, Neil Cornish and Scott Hughes for comments. This work has been partially supported thanks to a CSC scholarship.

### v2

@@ III.1 XMRI forest
Figure 2 shows the distribution of XMRI sources in the eccentricity–SNR plane, constructed from a randomly extracted sample consistent with the estimates of [5]. Most detectable sources cluster at low SNRs $\mathrm{SNR}<100$ and moderate to high eccentricities $e\sim 0.6-0.8$. Figure 3 further presents the relationship between SNR and the time to merger $T_{\mathrm{merge}}$ for the XMRI population, comparing 1-year (blue points) and 10-year (orange points) observation periods. As expected, longer observation times significantly enhance the accumulated SNR, especially for systems with shorter $T_{\mathrm{merge}}$.
@@
The results of our XMRI forest analysis are presented in Figure 4, along with several representative gravitational wave sources expected in the LISA band. Representative signals from massive black hole binaries , Galactic verification binaries, and characteristic EMRI harmonics are overlaid for comparison, together with the LISA instrumental sensitivity curve, consistent with those shown in Figure (1) of [1]. The yellow dots represent the superposition of individual XMRI strains binned in frequency, while the blue curve traces the non-zero strain values, illustrating the overall spectral shape of the forest. Notably, the XMRI background partially overlaps with the characteristic harmonics of polychromatic EMRI signals around frequencies of $\sim 1\,\mathrm{mHz}$, potentially introducing additional confusion noise in this region. Nevertheless, due to the low companion masses in XMRI systems, the background level remains relatively low and is not expected to substantially impede the detection of primary sources such as MBHBs or Galactic binaries.
@@ III.2 E-EMRI forest
Figure 7 shows the forest contribution from E-EMRI sources in our mixed catalog. These sources exhibit high strain amplitudes, producing a pronounced forest in the LISA band. The visible harmonic structure arises from their eccentric orbits, which generate gravitational radiation bursts across multiple harmonics. For the 10-year observation (lower panel), the improved frequency resolution leads to a finer and smoother spectral structure, particularly at higher frequencies $f\gtrsim 1\,\mathrm{mHz}$. Compared to the 1-year case, the extended observation reduces confusion at low frequencies and sharpens the forest contribution at high frequencies, in particular in the region where we expect polychromatic EMRIs to populate phase-space.
@@
The distribution of individual EMRI source strain amplitudes (grey dots) exhibits a distinct bifurcation, forming two visibly separated branches. This separation arises from the variation in source SNR. The lower branch is predominantly composed of sources with intermediate SNRs, whereas the upper branch originates from a small subset of loud sources. Owing to our background construction strategy—which maximizes spectral coverage by selecting realizations with the densest frequency occupation—the upper envelope is shaped by extreme cases, including sources reaching high SNRs, which contribute to the strain spectrum at high frequencies, leading to a visibly elevated branch relative to the bulk of the population.
@@ III.3 Extragalactic forests: The Andromeda galaxy
Figure 8 shows characteristic strain spectra for both E-EMRI and XMRI populations from M32 over a 1-year observation period with a handful of sources, as an illustration. The upper panel demonstrates that despite the substantial distance, many E-EMRI signals remain above LISA’s sensitivity curve in the $10^{-4}$ to $10^{-3}$ Hz band, particularly those involving higher-mass compact objects on highly eccentric orbits. In contrast, the lower panel shows that standard XMRI signals experience more severe attenuation, with most falling below the detection threshold across the frequency spectrum, although with a clear impact on polychromatic EMRIs. Obviously, understanding the shape of the forest when taking into account all close-by galaxies is important, and this will be addressed in a separate work.
@@ IV Conclusions
Acknowldegments
We thank Leor Barack, Neil Cornish and Scott Hughes for comments. This work has been partially supported by the National Natural Science Foundation of China (Grant No. 12247101), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06), the Natural Science Foundation of Gansu Province (No. 22JR5RA389) and the “111 Center” under Grant No. B20063. The support provided by China Scholarship Council (CSC) during a visit of Shao-Dong Zhao to the Universitat Politècnica de València is acknowledged.